OpenStudy (anonymous):

What is the exact volume of a sphere that has a radius of 7.5 ft?

OpenStudy (lyrae):

The volume of a sphere is \[\large V =\frac{ 4 }{ 3 } \times \pi \times r^3\]where r is the radius.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

That means 4/3*3.14*7.5^3

OpenStudy (anonymous):

that should help you get the exact volume

OpenStudy (anonymous):

how do you solve?

OpenStudy (anonymous):

ok... do you have a calculator?

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

divide this: 4/3

OpenStudy (anonymous):

now multiply that answer by 3.14

OpenStudy (anonymous):

When you get that answer,multiply it by 7.5 cubed

OpenStudy (anonymous):

and you should get your answer

OpenStudy (anonymous):

yes i do have a calculater

OpenStudy (anonymous):

now do all the math

OpenStudy (anonymous):

how do you divdide 4/3??

OpenStudy (anonymous):

put it into the calculator exactly like that

OpenStudy (anonymous):

1.33

OpenStudy (anonymous):

ok, is that answer still in the calculator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, multiply it by 3.14

OpenStudy (anonymous):

4.186

OpenStudy (anonymous):

multiply that answer by 7.5^3

OpenStudy (anonymous):

@kmachelle

OpenStudy (anonymous):

that wouldn't be correct

OpenStudy (anonymous):

i couldn't get it to work

OpenStudy (anonymous):

get what to work?

OpenStudy (anonymous):

the calculator wouldn't let me enter 7.5^3

OpenStudy (anonymous):

go to a calculator on google

OpenStudy (anonymous):

┌( ಠ_ಠ)┘

OpenStudy (anonymous):

k

OpenStudy (anonymous):

1765.96

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

is this a multiple choice question or no?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

@kmachelle

OpenStudy (anonymous):

check your work

OpenStudy (anonymous):

what did i get something wrong?

OpenStudy (anonymous):

@kmachelle

OpenStudy (anonymous):

when I did the whole equation I got a different answer

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

that's good but you do not need to simplify the number once you get it

OpenStudy (anonymous):

every time you put in a number keep it there the multiply the next number

OpenStudy (anonymous):

After

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

@kmachelle

OpenStudy (anonymous):

1766.25